### v1

# TSPDiffuser: Diffusion Models as Learned Samplers for Traveling Salesperson Path Planning Problems

Abstract
This paper presents TSPDiffuser, a novel data-driven path planner for traveling salesperson path planning problems (TSPPPs) in environments rich with obstacles. Given a set of destinations within obstacle maps, our objective is to efficiently find the shortest possible collision-free path that visits all the destinations. In TSPDiffuser, we train a diffusion model on a large collection of TSPPP instances and their respective solutions to generate plausible paths for unseen problem instances. The model can then be employed as a learned sampler to construct a roadmap that contains potential solutions with a small number of nodes and edges. This approach enables efficient and accurate estimation of traveling costs between destinations, effectively addressing the primary computational challenge in solving TSPPPs. Experimental evaluations with diverse synthetic and real-world indoor/outdoor environments demonstrate the effectiveness of TSPDiffuser over existing methods in terms of the trade-off between solution quality and computational time requirements.
Keywords: Path Planning, Diffusion Model, Traveling Salesperson Problem
1 Introduction
Consider an automated delivery robot tasked with delivering packages to several desks within a cluttered office. To ensure high levels of efficiency and safety, the robot must accomplish two tasks: determine the order of the desks to visit and find collision-free paths between consecutive destinations in obstacle-rich environments. Similar applications may include robotic security patrols [1] and inventory management [2], which we envision will be realized in the near future in unconstrained street corners and stores.
We are interested in a composite problem of traveling salesperson problems (TSPs) [3] and path planning introduced above, which we refer to as traveling salesperson path planning problems (TSPPPs). To solve a TSP for the shortest possible loop that visits all destinations, traveling costs must be estimated in advance for all pairs of destinations via path planning. While various implementations of fast TSP solvers are readily available (e.g., Google OR-Tools Routing Library [4]), path planning algorithms require longer computation times for higher solution quality (i.e., success rate and path optimality). This poses the main computational challenge in solving TSPPPs.
Machine learning has attracted increasing attention as a promising tool for enabling efficient, high-quality path planning in a data-driven manner. Given a large collection of path planning problem instances and their solutions, data-driven path planners learn a neural network to directly generate solution paths [5, 6, 7, 8, 9, 10, 11] or roadmaps/trees on which solution paths can be found efficiently via sampling-based path planning [12, 13, 14, 15, 16]. However, most existing data-driven planners are specialized for a single-query setup (i.e., require the exploration of the same environment for every pair of starts and goals), making them an inefficient choice for TSPPPs that need to execute path planning as many times as the square of the number of destinations.
In this work, we propose TSPDiffuser, a novel data-driven path planner designed to solve TSPPPs. As depicted in Fig. 1, we learn a diffusion model [17, 18] to generate plausible loop paths from obstacle and destination maps. Rather than attempting to generate a single solution path directly from a learned model, TSPDiffuser samples multiple paths to construct a roadmap for a higher success rate and path optimality. Because this roadmap highly likely contains potential solutions with a small number of nodes and edges, it alone can be used to efficiently estimate traveling costs for all destination pairs. In this way, TSPDiffuser aims to achieve both high solution quality and computational efficiency, addressing the main challenge for TSPPPs.
We extensively evaluate our approach using three synthetic environments with diverse obstacle layouts, as well as 30 outdoor and 15 indoor environments with real-world complexity [19, 20, 21]. Experimental results demonstrate that TSPDiffuser significantly improves the trade-off between solution quality and computation time over existing methods [10, 14, 22].
Figure 1: Solving TSPPPs with TSPDiffuser. (a) Problem instance. (b) Generating multiple paths (each shown in different colors) using a learned diffusion model. (c) Connecting paths to construct a roadmap. (d) Perform path planning and solving TSP on the roadmap to obtain a solution path.
2 Related Work
Path Planning and Data-Driven Planners
Path planning is a classical problem in AI and robotics, where the goal is to find the lowest-cost feasible path between two states (e.g., positions, poses). Particularly in continuous or high-dimensional state spaces, sampling-based approaches such as probabilistic roadmaps (PRM) [23], rapidly-exploring random trees (RRT) [24], and their asymptotically optimal versions [22] are often used. Recent work aims to leverage machine learning techniques to develop data-driven planners in various ways: learning neural networks that can directly generate paths [5, 6], learning sampling distributions for constructing roadmaps [12, 14, 15], or learning policies and/or value functions for tree-based search [13, 16]. Machine learning modules in such planners are often integrated into classical planning frameworks to ensure high success rates, thus inheriting the trade-off between planning quality and efficiency; the more search costs we spend, the more likely we are to obtain a feasible and optimal solution if one exists. Moreover, most of the existing data-driven planners except [14] consider a single-query setup and need to explore the environment for every pair of sources and destinations.
Traveling Salesperson Problem (TSP)
The traveling salesperson problem (TSP) is a combinatorial optimization problem that aims to find the shortest possible route visiting all given destinations exactly once before returning to the origin, given a set of destinations and pairwise traveling costs among them. TSP has a long history in theoretical computer science and operations research and has been extended to robotics tasks such as order picking in warehouses and vehicle routing [3]. Various forms of TSPs have been studied with practical assumptions on agents (e.g., vehicles, unmanned aerial vehicles, or teams of such moving agents), including Dubins TSP [25], stochastic TSP [26], variable-speed TSP [27], trajectory-based TSP [28], and vehicle routing problems [29]. The proposed TSPPP can be viewed as a specific case of metric TSPs, where traveling costs between destinations are given by the length of shortest possible collision-free paths and thus only satisfy the triangle inequality. Studies in this direction have been relatively limited [30, 31], and no prior work has considered its path planning aspect, i.e., how to accurately and efficiently evaluate traveling costs in obstacle-rich environments.
Diffusion Models
Diffusion models belong to a family of probabilistic generative models, which are capable of generating high-quality data through an iterative denoising process. This is achieved by progressively injecting noise to disrupt input data, and learning the reverse process that gradually reconstructs the original data back from complete noise (see [18] for an introductory survey.) Diffusion models have led to significant advances especially in image generation [32, 33], and have also begun to be used in robot learning for modeling environmental dynamics [10, 11]. A recent work has attempted to solve Euclidean TSPs using diffusion models [34].
3 Preliminaries
3.1 Problem Definition
Fig. 1 provides an overview of our work. An instance for TSPPPs is represented by an environmental map with obstacles and a set of destinations (Fig. 1a). The solution is expected to be a path that visits all the destinations while avoiding obstacles and returning to the starting point (Fig. 1d). The quality of solution paths is measured by feasibility (i.e., if the path is collision-free) and optimality (i.e., how close the path is to the shortest one.) The performance of planners is evaluated by the trade-off between solution quality and computational efficiency, or in practice, the computation time required to derive solutions.
Our primary objective is to improve this trade-off for solving TSPPPs. Particularly in mobile robotics applications, scalability with respect to the number of destinations can inherently be limited due to battery constraints and the variability of environments. Instead, quick and high-quality path planning becomes more critical to increase the overall throughput in task execution, as it is used thoroughly to estimate traveling costs between destinations.
Concretely, let $\Omega=[-1,1]^{2}$ be a continuous state space, where the free (i.e., non-obstacle) space is given by $\Omega_{\mathrm{free}}\subset\Omega$ and the position of a robotic agent, as well as its destinations, are represented by a 2D real-valued vector $p\in\Omega_{\mathrm{free}}$. We represent the free space by a roadmap $G=(V,E)$, where nodes $V\subset\Omega_{\mathrm{free}}$ represent positions and edges $E\subset V\times V$ connect the nodes with their neighbors if there are no obstacles between them. Valid edges have the Euclidean distance between the positions as edge costs. Let $\Omega_{\mathrm{dst}}\subset\Omega_{\mathrm{free}}$ be a set of destinations to visit, which we assume to be included in the vertex set as well, i.e., $\Omega_{\mathrm{dst}}\subset V$. With a graph search algorithm, we can find a shortest path and its cost between destinations on the constructed roadmap.
3.2 Denoising Diffusion Probabilistic Models (DDPMs)
We propose to leverage diffusion models for constructing a roadmap with a small number of significant nodes and edges. We will particularly adopt denoising diffusion probabilistic models (DDPMs) [17] that model perturbation and denoising processes by Markov chains, although any diffusion model could be used in principle. Let $\tau^{0}$ be a clean, original data sample and $\tau^{i}$ be its perturbed version by progressively injecting noise $i$ times ($1\leq i\leq I$). In DDPMs, the reverse transition from $\tau^{i}$ to $\tau^{i-1}$ is typically given by the Gaussian with learnable parameters $\theta$ [10]:
$p_{\theta}(\tau^{i-1}\mid\tau^{i})=\mathcal{N}(\tau^{i-1}\mid\mu_{\theta}(\tau^{i},i),\Sigma^{i}),$ (1)
where $\Sigma^{i}$ is a covariance matrix. The function $\mu_{\theta}(\tau^{i},i)$ is parameterized by noise estimator $\varepsilon_{\theta}$:
$\mu_{\theta}(\tau^{i},i)=\frac{1}{\sqrt{\alpha^{i}}}\left(\tau^{i}-\frac{\beta^{i}}{\sqrt{1-\bar{\alpha}^{i}}}\varepsilon_{\theta}(\tau^{i},i)\right),$ (2)
where $\alpha^{i}$, $\bar{\alpha}^{i}$ and $\beta^{i}$ are inter-dependent hyper-parameters chosen appropriately. The function $\varepsilon_{\theta}(\tau^{i},i)$ is given by neural networks such as U-Net [35], and is learned from training data and their perturbed versions (see [17] for the complete introduction.)
Learned DDPMs can generate clean data from complete noise $\mathcal{N}(\mathbf{0},\mathbf{I})$ by repeatedly applying the denoising kernel in Eq. (1). Particularly for trajectory generation [10, 11], reward (value)-guided sampling encourages generated samples to meet an objective represented by a differentiable reward (or value) function $\mathcal{J}(\tau)$. Specifically, for sample $\tau^{i}$, the denoising step is extended as follows [10]:
$\tau^{i-1}\sim\mathcal{N}(\mu+\alpha\Sigma\nabla\mathcal{J}(\mu),\Sigma^{i}),$ (3)
where $\mu=\mu_{\theta}(\tau^{i},i)$ and $\alpha$ is a scale parameter.
Figure 2: Model Architecture. We represent a TSPPP instance using binary and distance images of obstacles and destinations, which are stacked and encoded into a unified feature map. Noisy position data, concatenated with denoising timestep information and the feature vector extracted from the feature map, are fed into a diffusion model to reconstruct the original clean path.
4 TSPDiffuser
In TSPDiffuser, we extend a diffusion model to generate plausible loop paths for a given TSPPP instance (Fig. 1b). The generated paths are interconnected to form a roadmap with a small number of vital nodes and edges around potential solutions (Fig. 1c). This efficient construction of a high-quality roadmap contributes to the overall throughput improvement. By carrying out path planning and solving the TSP on this roadmap, we derive the solution path (Fig. 1d).
4.1 Model Architecture
Fig. 2 illustrates the proposed model architecture. We define a sample $\tau^{i}$ by a path, i.e., a sequence of positions $\tau^{i}=(p^{i}_{0},\dots,p^{i}_{T}),\;p^{i}_{t}\in\Omega$, that tours all destinations while avoiding obstacles. Because input instances provide the layout of obstacles and destinations, we propose to feed this information into the noise estimation network $\varepsilon_{\theta}$, effectively conditioning its generation results.
Specifically, we represent obstacles and destinations by binary images and their distance images obtained via distance transform. These images are stacked and encoded into a $d$-dimensional unified feature map $F$ using fully-convolutional networks. The feature vector extracted at location $p$ from the feature map, denoted as $F(p)\in\mathbb{R}^{d}$, describes a local context about whether there are obstacles and destinations nearby. Our noise estimator takes the form of $\varepsilon_{\theta}(\tau^{i},i,F(\tau^{i}))$, where $F(\tau^{i})=(F(p^{i}_{1}),\dots,F(p^{i}_{T}))$ is a sequence of feature vectors extracted along $\tau^{i}=(p^{i}_{1},\dots,p^{i}_{T})$. In practical implementations, we concatenate the position $p^{i}_{t}$, the denoising timestep $i$ encoded into a vector via Fourier transform, and the feature vector $F(p^{i}_{t})$ to form the input to the noise estimation network.
4.2 Constructing Roadmaps using Learned Diffusion Models
Despite their high generation capability, we observe two limitations for diffusion models when used in path planning. Firstly, sample generation by diffusion models is typically computationally intensive, which is not necessarily suitable for practical applications where higher throughput with limited computation resources is appreciated. Although many extensions have been developed to mitigate this problem (e.g., [36, 37]), efficient sampling often sacrifices the quality of generated samples. Secondly, even with reward-guided sampling, there is no guarantee that generated paths satisfy constraints imposed in TSPPPs, i.e., visiting all destinations in a single loop, while avoiding collisions with any single obstacle. This makes it hard to solve TSPPPs by diffusion models alone.
Addressing these limitations, we propose to leverage a diffusion model as a learned sampler for constructing a roadmap $G=(V,E)$. We sample multiple loop paths from a learned diffusion model with relatively low timestep granularity (e.g., $I=5$, as shown in Fig. 1b), and interconnect these paths and destinations to form a roadmap (Fig. 1c). The resulting roadmap contains potential solution paths with a small number of important nodes and edges. This approach is more accurate than attempting to generate a single high-quality path that satisfies the TSPPP constraints, and is more efficient than building a conventional roadmap that covers the entire free space in the environment.
Alg. 1 describes the proposed algorithm. Let $\{\tau_{1},\dots,\tau_{M}\}$ be a set of paths generated from a learned diffusion model, where $\tau_{m}=(p_{m,1},\dots,p_{m,T_{m}})$ is a sequence of positions. All the positions contained in the generated paths are regarded as nodes $V$ (L3). With a validity check function $\texttt{isvalid}(\cdot,\cdot;\Omega_{\mathrm{free}})$ that returns true if there are no obstacles between a pair of positions, we first connect nodes along each generated path to define edges $E$ (L4-L7). After adding destinations $\Omega_{\mathrm{dst}}$ and nodes around obstacle boundaries $\{b_{1},\dots,b_{N}\}$ to $V$ (L9), we further span edges between each node and its $K$ nearest neighbors to $E$ (L10-L15).
Algorithm 1 Construct Roadmaps with Diffusion Models
1:A collection of paths $\{\tau_{1},\dots,\tau_{M}\}$, $\tau_{m}=(p_{m,1},\dots,p_{m,T_{m}})$, generated from a learned diffusion model; a set of destinations $\Omega_{\mathrm{dst}}$; validity check function $\texttt{isvalid}(\cdot,\cdot;\Omega_{\mathrm{free}})$
2:Roadmap $G=(V,E)$
3:$V,E\leftarrow\{\},\{\}$
4:for $m=1,\dots,M$ do $\triangleright$ Connect nodes along paths
5:     $V\leftarrow V\cup\{p_{m,t}\mid p_{m,t}\in\tau_{m}\}$
6:     for $t\in 1,\dots,T_{m}-1$ do
7:         if $\texttt{isvalid}(p_{m,t},p_{m,t+1};\Omega_{\mathrm{free}})$ then $E=E\cup\{(v,w)\}$
8:         end if
9:     end for
10:end for
11:$V\leftarrow V\cup\Omega_{\mathrm{dst}}\cup\{b_{1},\dots,b_{N}\}$triangleright Add destinations and $N$ boundary nodes
12:for $v\in V$ do $\triangleright$ Connect nodes with their neighbors
13:     for $w\in\texttt{K-NN}(v)$ do $\triangleright$ $K$ neighbors around $v$
14:         if $\texttt{isvalid}(v,w;\Omega_{\mathrm{free}})$ then $E=E\cup\{(v,w)\}$
15:         end if
16:     end for
17:end for
4.3 Solving TSPPP
On the constructed roadmap, we perform path planning to calculate the shortest-path distance as traveling costs between each pair of destinations. We then solve a TSP using the calculated traveling costs. This gives us the order of destinations to visit with the total path cost minimized among all possible tours. Finally, we stitch the shortest paths between destinations in the determined order to generate the solution path for TSPPPs.
5 Experiments
5.1 Planning Environments
We extensively evaluate TSPDiffuser using both synthetic and real-world indoor/outdoor navigation environments, each with various layouts of obstacles and destinations on 128x128 sized maps:
  • Synthetic: Three different synthetic environments with different obstacle layouts; Standard with twenty random 20x20 square-shaped obstacles, More Obstacles with forty random 10x10 square-shaped obstacles, and Larger Obstacles with ten random 40x40 square-shaped obstacles.
  • City/Street Map [19]: Random areas cropped from 30 city/street maps such as Berlin, London, and New York.
  • Indoor Navigation Dataset [20]: Random areas cropped from 15 indoor office/living environments for mobile robot navigation tasks, which was generated from Gibson Database [21].
For each environment, we created 100 instances by randomizing the locations of obstacles (for Standard, More Obstacle, and Larger Obstacle) and cropped areas (for City/Street Map and Indoor Navigation Dataset). For each instance, ten random destinations were sampled from the largest non-obstacle area. The ground-truth, i.e., (pseudo-) optimal, path for each instance was obtained by computing traveling costs among destinations using PRM [22] with 3,000 nodes and solving the TSP using or-tools [4].
In addition, we generated approximately 100,000 instances of the Standard environment as a training dataset for TSPDiffuser and other data-driven baselines presented later. This verifies the generalization ability of these planners, especially from synthetic to real-world environments. Ground-truth paths in the training dataset were repeated until the total number of temporal steps exceeded 256, and then trimmed to have a fixed step of 256. This process allowed us to easily stack multiple ground-truth paths to form a mini-batch during model training.
5.2 Evaluation Setups
Evaluation Metric
For each environment, we calculated the success rate weighted by path length (SPL) [38] $\frac{1}{100}\sum_{i=1}^{100}S_{i}\frac{\hat{l}_{i}}{\max(l_{i},\hat{l}_{i})}$ for 100 instances, where $S_{i}$ is the binary indicator that takes $S_{i}=1$ if the method found a solution path for the $i$-th instance and $S_{i}=0$ otherwise, and $l_{i},\hat{l}_{i}$ are the lengths of the solution and ground-truth paths. We evaluate the trend of SPL scores while changing hyper-parameters of a method that affect the trade-off between the solution quality and required computation times, such as the number of nodes for PRMs and that of generated paths $M$ for TSPDiffuser. All the evaluation procedures were run on the Intel® Xeon® CPU @ 2.20GHz.
Baselines
Due to the absence of prior work that solves the same problem, we have extended existing path planning algorithms for baseline methods. All the baselines below are able to construct a roadmap to evaluate traveling costs for all combinations of destinations (i.e., support multi-query setups), which is then followed by the same solution procedure shown in Sec. 4.3.
  • Probabilistic Roadmap (PRM) [22]: A classical sampling-based path planning algorithm that is also used to generate ground-truth paths. Roadmaps were constructed by sampling $M_{\mathrm{prm}}\in\{200,500,1000,2000\}$ nodes uniformly from non-obstacle regions, and then connecting the nodes based on $r$-neighbor search, where $r$ was determined adaptively based on [22].
  • Critical PRM (CPRM) [14]: A data-driven path planner that learns to detect ‘critical’ regions in environments and prioritizes sampling from those regions for constructing a roadmap. In this baseline, we directly regard regions around ground-truth paths as critical, and learn a U-Net to predict those regions. Sampled nodes are connected based on $r$-neighbors, where $r$ was determined in the same way as that of PRM. [^1] Although the original CPRM connects non-critical nodes to all critical nodes, we employed $r$-neighbors because such full connections required a long time despite limited solution quality in our problem setup. The number of nodes was selected among $M_{\mathrm{cprm}}\in\{100,200,300,400\}$ to see how SPL scores and required times can change.
  • Diffuser [10]: A state-of-the-art trajectory generation method using diffusion models. For a fair comparison, we employed the same diffusion model architecture as that of TSPDiffuser in Sec. 4.1. A single loop path is generated from a learned model and then connected with destinations and boundary nodes to construct a roadmap. This baseline can also be viewed as a degraded version of TSPDiffuser where the number of paths is limited to $M=1$.
Implementation Details
TSPDiffuser consists of the problem instance encoder network and the noise estimation network. As the instance encoder, we used a U-Net [35] with four encoding and decoding layers from the VGG backbone [39], which outputs $d=256$-dimensional feature maps. The noise estimation network was a 1D U-Net with residual blocks and self-attention modules, where the input is the sequence of concatenations of positions, denoising timesteps encoded into 16-dimensional real-valued vectors via Fourier transform, and the 256-dimensional feature vectors extracted from the feature maps. These two modules were trained end-to-end with the standard DDPM training pipeline [17] for 300 epochs with a batch size of 32, using the AdamW optimizer [40] with the cosine annealing learning rate scheduler [41]. The granularity of denoising timesteps, i.e., $I$ in Sec. 3.2, was set to 500 for training, and 5 for sampling. For roadmap construction, the number of sampled paths $M$ were set to either one of $M\in\{5,10,20,30\}$, where larger $M$ will result in higher SPL scores with more computation times. The reward guidance, i.e., $\nabla\mathcal{J}(\tau)$ in Eq. (3), is given by the sum of distance images from obstacles and destinations, which guides output positions to avoid obstacles and approach destinations. The number of neighbors considered when spanning edges among paths, i.e., $K$ in Alg. 1, were set to $K=5$ unless specified otherwise. The number of boundary nodes, i.e., $N$ in Alg. 1, were set adaptively by detecting boundary pixels of obstacles.
5.3 Results
Quantitative Results
Fig. 3a presents the relationships between SPL scores and computation times for all the methods. Despite being trained on synthetic data, TSPDiffuser effectively generalizes to real-world complex environments (a4 and a5), strongly indicating its transferability to real robots. It achieves high SPL scores with processing speeds that are almost 2x to 4x faster compared to the other baselines. Diffuser operates faster than TSPDiffuser, but its SPL score is significantly lower because a single sampling by diffusion models does not necessarily guarantee feasible or near-optimal paths. CPRM achieved comparable performance with much fewer nodes compared to PRM. Fig. 3b further elaborates on the computation times of TSPDiffuser and PRM. The construction of the roadmap is the main computational bottleneck. TSPDiffuser effectively tackles this problem by limiting the nodes and edges to be sampled around potential solution paths.
Figure 3: Quantitative Results. (a1-a5, d1-d3) SPL scores and required computation times. (b) Specific time breakdown for TSPDiffuser and PRM. (c) Ablation study.
Ablation Study
We evaluated variants of TSPDiffuser with different hyper-parameters: (1) the number of $K$-neighbors to connect edges from $K\in\{3,5,9,15\}$, (2) the granularity of denoising timesteps from $I\in\{5,20,50,100\}$, and (3) whether or not nodes along obstacle boundaries are added to roadmaps, to see how these choices impact SPL scores and required times. As summarized in Fig. 3c, adding boundary nodes is crucial. The number of $K$-neighbors is also important to ensure high SPL scores. Despite increased computation times, no significant improvements were confirmed for higher timestep granularity $I\geq 20$.
Results for More Destinations
We further investigated the scalability of TSPDiffuser with respect to the number of destinations: 20, 50, and 100, on the Standard environment. Based on the ablation study, we evaluated TSPDiffuser with $K=15$ to pursue more optimal solutions at the cost of required computation times. As summarized in Fig. 3d, TSPDiffuser ($K=15$) consistently outperforms the other baselines. It can find solutions with an SPL higher than $0.98$ in around 10 seconds for instances with 100 destinations, whereas PRM requires nearly 50 seconds to achieve the same level of performance.
Qualitative Results
Fig. 4 visualizes typical solution paths and roadmaps obtained by each method. Compared to PRM, which uniformly samples nodes, data-driven planners such as CPRM, Diffuser, and TSPDiffuser concentrate more nodes around ground-truth paths. Diffuser sometimes failed to find feasible paths connecting all destinations without collisions. TSPDiffuser can find collision-free paths more reliably by generating multiple diverse paths to form a roadmap.
Figure 4: Qualitative Results. Roadmaps and solution paths are visualized with gray and black lines (if solved). Obstacle regions are colored in blue, and destinations are marked with red circles.
Figure 5: Failure case.
5.4 Limitations
Fig. 5 shows a simple case for a failure mode of TSPDiffuser. The diffusion model generates straight paths between two destinations despite the presence of obstacles in between. We speculate that this is partly due to the absence of highly skewed loop paths (as indicated by the black line) in the training data. Moreover, the current version of TSPDiffuser only supports the 2D state space, whereas practical robotics applications can sometimes involve more complex state spaces [14]. This requires extensions to the architectures for the instance encoder and noise estimation networks, which we seek in future work.
6 Conclusion
We introduced a novel data-driven path planner named TSPDiffuser for traveling salesperson path planning problems (TSPPP) in obstacle-rich environments. Given a pair of obstacle and destination maps, TSPDiffuser samples plausible paths from a learned diffusion model and connects them to form a roadmap with a small number of essential nodes and edges. This improves the balance between solution quality and efficiency in solving TSPPPs. Our evaluation has confirmed the effectiveness of TSPDiffuser over existing classical and state-of-the-art data-driven path planners in both synthetic and real-world indoor/outdoor navigation environments.